OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (triciaal):

point slope formula for 2 points on a line (x, y) any point and (2, -4) given point (x1, y1) slope m = 3/5 y-y1= m(x - x1) (y--4 )= 3/5(x -2) y + 4 =3/5x -3/5*2 sure you can finish to y =